OpenStudy (anonymous):

iis the sequence 1,3,5,7 geometric? is so find r

OpenStudy (anonymous):

no its not geometric - its arithematic It goes by adding 2 to the next item

OpenStudy (anonymous):

7- 5 = 2 5-3 = 2 3- 1 = 2 so d = 2